If the cost price of 15 pens is equal to the selling price of 20 pens , find the loss per cent

Grade:7

2 Answers

sonika p
6 years ago
let, cost of 1 pen = x
cost of 15pens =15x
selling price = cost of 20pens =20x
loss percentage = ((cost price-selling price)/cost price)*100
                          = ((15x-20x)/15x=(-1/3)*100
                          =33.33%
thus loss is 33.33 percent
Nitesh
6 years ago
Given cost price of 15 pen= selling price of 20 penLet cost price of 15 pen =X So cost price of one pen= X/15Selling price of 20 pens = X So selling price of 1 pen = X/20Loss on per pen=cost price- sell price=》X/15 -X/20=X/60 Now loss percent=loss×100/cost priceLoss percent=X/60 ×100 × 15/X=》 LOSS PERCENT=25. hence the answer is 25°/•